What is the 2001st number in the sequence: 2, 5, 8, 11,...?
Found 2 solutions by dyakobovitch, Alan3354:
Answer by dyakobovitch(40)

You are dealing with an arithemetic sequence with d=r, a(sub1)=2, and n=2001.

Your formula for finding the a(subn) term of an arithmetic sequence is
a(subn)=a(sub1) + (n-1)d = 2 + (2001-1)(3) = 6002.

I hope this help! Feel free to leave comments or concerns. -- David

Answer by Alan3354(69443)

What is the 2001st number in the sequence: 2, 5, 8, 11,...?
r = 3
-------------
If 2 is the 1st:
n1 = 2
n2 = 5 (= 2 + 3*(n-1))
n3 = 8 (= 2 + 4*(n-1))
.
.
.
n2001 = 2 + (2001-1)*r
n2001 = 6002
--------------------
If 2 is the 0th, then it's 6005